Find the equation of the line using the point-slope formula. Write the final equation using the slope-intercept form.
perpendicular to
3y = x − 4
and passes through the point
(−2, 1)

c

Answer by mananth(16946)

3 y = 1 x -4
Divide by 3
y = 1/ 3 x + -1 1/3
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
Because m1*m2 =-1
The slope of the required line will be -3

m= -3 ,point ( -2 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = 6 + b
b= -5
m= -3
The required equation is y = -3 x -5
m.ananth@hotmail.ca